OpenStudy (anonymous):

Zach borrowed four thousand dollars from his grandfather to buy a used car so he could travel to and from work. After working for 3 years, he has saved five thousand nine hundred twenty dollars to pay back to his grandfather. At what interest rate did Zach borrow the money?

OpenStudy (apoorvk):

i am assuming simple interest as you havent mentioned anything now principal amount=P=4000 rate of interest per annum=R=? t=no. of years borrowed for=3 now he is paying an interest of "I'=total final amount-borrowed principal= 1900 now I=(P x R x t)/100 1900=(4000x R x 3)/100 solve, you should get the rate of interest per annum. multiply by 100 to get interest %age.

OpenStudy (lifeisadangerousgame):

I agree.^

OpenStudy (anonymous):

lol life

OpenStudy (anonymous):

what does that equal btw

OpenStudy (lifeisadangerousgame):

Whatever it should equal.

OpenStudy (apoorvk):

LOL

OpenStudy (anonymous):

i need help again

OpenStudy (lifeisadangerousgame):

Well it's true!

OpenStudy (anonymous):

with? what@jack

OpenStudy (anonymous):

understanding this guys answer...its right but what is the percentage

OpenStudy (anonymous):

ooh a medal...thanks@life and jack

OpenStudy (lifeisadangerousgame):

Cuz I'm awesome like that...Jack you need to figure out the answer Lazy Doctors now! Humph. I'll never go into surgery with you around! You might be too lazy to figure out if I need my lungs!

OpenStudy (anonymous):

HAHAHAHAHAhah

OpenStudy (anonymous):

u dont need your lungs id be happy to remove them heh heh heh

OpenStudy (lifeisadangerousgame):

I'm pretty sure I need all 4 humph.

OpenStudy (anonymous):

werewolf trash

OpenStudy (anonymous):

LMAO

OpenStudy (lifeisadangerousgame):

Dog poop!

OpenStudy (anonymous):

hairball

OpenStudy (lifeisadangerousgame):

Buttbreath!

OpenStudy (anonymous):

WOW

OpenStudy (anonymous):

buttbreath? thats pretty serious

OpenStudy (anonymous):

your the dog not me

OpenStudy (anonymous):

thats biskerz your lookin at

OpenStudy (lifeisadangerousgame):

YOU JUTS SAID HAIR BALL! CATS get hairballs!

OpenStudy (anonymous):

werewolfs get em too

OpenStudy (lifeisadangerousgame):

So does your dog HAHAHA

OpenStudy (anonymous):

no he doesnt take it back

OpenStudy (lifeisadangerousgame):

NEVERRR

OpenStudy (anonymous):

i hate u your not my fan anymore u insulted my cute dog

OpenStudy (lifeisadangerousgame):

You insulted my werewolf lineage!

OpenStudy (anonymous):

im sorry

OpenStudy (lifeisadangerousgame):

ME too cabbagebrain..

OpenStudy (anonymous):

what ever

OpenStudy (lifeisadangerousgame):

I have several bees nest in my yard.

OpenStudy (anonymous):

i gtg

OpenStudy (anonymous):

bye jack

OpenStudy (lifeisadangerousgame):

bye!

OpenStudy (anonymous):

no more fighting now

OpenStudy (anonymous):

Simple Interest formula: I = p r t <=> r = I/ p * t The difference amount between borrowed and paid back is the interest: I = $5,920 - $4,000 = $1,920 => Interest rate: r = 1,920 / 4,000 * 3 = .16 = 16%

OpenStudy (anonymous):

i remember this from chat almost a month ago

OpenStudy (anonymous):

XD that was amazing, and whats even better at the end the answer was there "16%"